OpenStudy (anonymous):

How do I solve the integral of (10x^2)/((x+1)(x^2+1))?

OpenStudy (jamesj):

You need to break it up into partial fractions.

sam (.sam.):

\[\frac{A}{x+1}+\frac{Bx+C}{x ^{2}+1}\]

OpenStudy (anonymous):

I am getting stuck on how to solve once I have broken it up, specifically. I tried solving by setting the numerator equal to the original numerator \[10x^2 = A(x^2+1) + (Bx+C)(x+1)\] But all the coefficients are 0 other than 10x squared

OpenStudy (jamesj):

Yes, what you have is a system of linear equations in A, B and C. But not all of A, B and C are zero.

OpenStudy (anonymous):

Essentially I am getting confused because to evaluate for x^0, I get A+B+C, for x^1 I get just C, and for x^2 I get A+B. Since C has to be zero, how do I accommodate to make A+B+C zero?

OpenStudy (jamesj):

Your equation above is \[ x^2=A(x^2+1)+(Bx+C)(x+1) \] \[ = (A+B)x^2 + (B + C)x + (A+C) \] Hence A + B = 1 B + C = 0 A + C = 0

OpenStudy (anonymous):

Oh wow, I can't believe I missed something so simple :P Thank you very much